M and N are the midpoints of side AB and AC of triangle ABC
1.If MN=2 x-5 and BC = X + 2, find x
2. If angle AMN=(2y + 25) and Angle B = (3Y - 5), find Y and Angle B

Grade:9

3 Answers

Vikas TU
4 years ago
Dear student 
According to midpoint theorem 
MN = ½ BC 
So, 2x -5 = ½ (x+2) 
4x – 10 = x+ 2 
3x = 12 
x = 4 
ii). Angle AMN is appx equal to B 
So , 2y +5 = 3y -5 
y = 10 
 
Vikas TU
4 years ago
Dear student 
Angle B would be equal to 
30 -5 = 25 degree 
Hope this helps ….....................
Good Luck 
Aditya Gupta
4 years ago
hello student, once again the answer by vikas for the 2nd part is wrong and immature.
we know by mid pt theorem that angle B= angle AMN.
so, 2y+25 = 3y – 5
or y= 30
now, B= 3y – 5
= 90 – 5
or B= 85
kindly approve :)